OpenStudy (mollykenney):

8r-5 > or equal 6r -1 or 10 + 3a < or equal -5

OpenStudy (aum):

\(8r-5 \ge6r -1 \) Add \(-6r\) to both sides: \(2r-5 \ge -1 \) Add \(5\) to both sides: \(2r \ge 4 \)

OpenStudy (aum):

Divide both sides by 2: \(r \ge 2\)

OpenStudy (aum):

\(10 + 3a \le -5\) Add \(-10\) to both sides: \(3a \le -15\) divide both sides by 3: \(a \le -5\)

OpenStudy (mollykenney):

Thank you so much